# Twist, splay, and uniform domains in ferroelectric nematic liquid crystals

### v1

@@ Abstract
The recently-discovered ferroelectric nematic ($\mathrm{N}_{\mathrm{F}}$) liquid crystal presents a host of defect phenomena due to its unique polar nature and long-ranged electrostatic interactions. Like the solid state ferroelectrics, the depolarization field in the material favors a spontaneous spatial variation of the polarization $\mathbf{P}$, manifesting in myriad ways including a twist in the bulk and different arrangements of alternating polarization domains. Unlike the solid-state ferroelectrics with a bulk crystalline structure, the configuration of the $\mathrm{N}_{\mathrm{F}}$ fluids is determined not only by the reduction of depolarization fields but also by the alignment of molecules at interfaces. In this work, we consider an $\mathrm{N}_{\mathrm{F}}$ confined to a thin cell, pre-patterned with various types of apolar surface anchoring produced by photoalignment. For uniform planar alignment, we find that the sample forms a series of striped domains. For a cell pre-patterned with a radial +1 defect pattern, the $\mathrm{N}_{\mathrm{F}}$ breaks up into “pie-slice” polarization domains. We calculate the elastic and electrostatic energy balance which determines the observed configurations and demonstrate that the electrostatic interactions tend to decrease the characteristic domain size $\lambda$ while the elastic and surface anchoring interactions facilitate a larger $\lambda$. We also demonstrate that ionic screening mitigates electrostatic interactions, increasing $\lambda$ and, above some critical concentration, eliminating the domains altogether.
I Introduction and Background
@@
The polar ordering of the recently discovered ferroelectric nematic ($\mathrm{N}_{\mathrm{F}}$) liquid crystal Mandle et al. (2017a); Nishikawa et al. (2017a); Sebastián et al. (2020); Chen et al. (2020) creates a fascinating interplay between the elasticity, surface interactions, and the electrostatic energy associated with spatial variations of the spontaneous polarization $\mathbf{P}$. The ensuing $\mathrm{N}_{\mathrm{F}}$ structures are not restricted by crystallographic axes and can be studied by polarizing optical microscopy since, in the materials explored so far, the polarization $\mathbf{P}$ is along the optic axis, or the director $\hat{\mathbf{n}}$.
@@
$\mathrm{N}_{\mathrm{F}}$ structures tend to avoid a splay deformation (divergence of $\mathbf{P}$) since it creates a bound charge of a bulk density $\rho=-\nabla\cdot\mathbf{P}$ and increases the electrostatic energy. Polarization-related charges are also avoided at surfaces and at domain walls. For example, a uniform polarization, $\mathbf{P}(x,y,z)=\text{const.}$., would deposit charges at the opposite ends of the sample and create a strong, energetically costly depolarization field: $E=-P/\epsilon\epsilon_{0}\approx 10^{8}~{}\mathrm{V}/\mathrm{m}^{2}$; here $\epsilon\approx(10-100)$ Adaka et al. (2024); Erkoreka and Martinez-Perdiguero (2024) and $P\approx(3-7)\times 10^{-2}~{}\mathrm{C}/\mathrm{m}^{2}$ are the typical permittivity and polarization of the $\mathrm{N}_{\mathrm{F}}$ phase. Polydomain textures of thin $\mathrm{N}_{\mathrm{F}}$ films that impose no preferred in-plane orientation of $\mathbf{P}$, such as a film supported by an isotropic fluid Basnet et al. (2022); Kumari et al. (2023) or freely suspended in air Hedlund et al. (2025), present clear evidence of these tendencies. First, the spontaneous polarization is everywhere in the plane of the film, avoiding surface charges, which would occur whenever $\mathbf{P}$ is tilted. Second, the in-plane textures are dominated by two types of domains, in which $\mathbf{P}$ is either uniform or bends into circular vortices, thus avoiding splay deformations and associated space charge in the bulk. Domain walls separating these domains adopt the shapes of conic sections, such as parabolas and hyperbolas Basnet et al. (2022); Kumari et al. (2023); Hedlund et al. (2025).
@@
The $\mathrm{N}_{\mathrm{F}}$ structure changes dramatically when one of the film’s surfaces imposes a unidirectional alignment of $\mathbf{P}$ and the other is azimuthally degenerate. In this case, one might expect a uniform state since circular vortices are not compatible with the unidirectional surface alignment. However, experiments Kumari et al. (2024a) demonstrate that instead of being uniform, the polarization twists around the film normal, so that the vectors $\mathbf{P}$ are antiparallel to each other at the bottom and the top surfaces, thus mitigating the depolarization effect. Since the $\mathrm{N}_{\mathrm{F}}$ molecules are not chiral, the film splits into equal-width stripes of alternating twist handedness. An analogous twisting phenomenon was predicted by Khachaturyan in 1975 Khachaturyan (1975) for an infinitely long cylindrical $\mathrm{N}_{\mathrm{F}}$ sample, in which $\mathbf{P}$ twists around the axis of the cylinder. In both experiments Kumari et al. (2024a) and in Khachaturyan’s model Khachaturyan (1975), the twisted structures arise from the balance of electrostatic and elastic energies. Clearly, this balance should be affected by the geometry of confinement (e.g., a film with a large lateral extension or a long cylinder), by surface anchoring at the bounding plates, and by the free ions capable of at least partial screening of bound charges.
@@
In this work, we explore experimentally and theoretically the interplay between electrostatic and elastic energies in the $\mathrm{N}_{\mathrm{F}}$ domain structures, taking into account the effects of surface anchoring and ionic screening. The surface anchoring in the experiments is designed to be in-plane apolar, or “bidirectional”, by using a photoalignment technique Guo et al. (2016a), Guo et al. (2016b). There are two reasons: First, apolar anchoring should avoid twists that are artificially created by the antiparallel assembly of two plates with a unidirectional alignment of $\mathbf{P}$, which happens for mechanically rubbed plates. Second, the photoalignment technique allows one to impose various patterns of surface alignment of $\mathbf{P}$ (including an unwelcome splay), thereby exploring the electrostatics-elasticity balance in different contexts.
@@
We find that, generally, the electrostatic interaction prefers a spatial modulation of $\mathbf{P}$ with a characteristic size $\lambda$ (e.g., the twist pitch or the domain size), generating an energetic contribution that increases with $\lambda$. Conversely, any such modulation incurs an elastic or anchoring energy penalty, which decreases with $\lambda$. Thus, the balance between elastic (or anchoring) energy and the electrostatic interaction generates a preferred value of $\lambda$, leading to the various patterns considered here. The two main motifs of these patterns are (i) neighboring domains with a uniform polarization that flips by $\pi$ when one moves from one domain to the next and (ii) domains with additional left- and right-handed $\pi$-twists of polarization along the axis orthogonal to the cell’s plane. As a rule, the first type of pattern occurs in thin (micron) films, while the second pattern is prevalent in thicker slabs.
Note that the nematic director $\hat{\mathbf{n}}$ in the $\mathrm{N}_{\mathrm{F}}$ does not have to align with the polarization $\mathbf{P}$ Sebastián et al. (2022). In this case, we have to consider the flexoelectric coupling (a term proportional to $-(\mathbf{P}\cdot\hat{\mathbf{n}})(\nabla\cdot\hat{\mathbf{n}})$ in the free energy). Such a term, along with a treatment which considers the nematic and polar order separately, can lead to other interesting phases and instabilities in these materials Paik and Selinger (2024); Sebastián et al. (2022); Vafa and Doostmohammadi (2025). Here, however, we will assume that $\mathbf{P}$ is always parallel to $\hat{\mathbf{n}}$ and $\mathbf{P}\cdot\hat{\mathbf{n}}=\pm P_{0}$, which occurs when the coupling is sufficiently strong. Our primary interest will be the effects of the uncompensated bound charges at sample boundaries and due to non-vanishing $\nabla\cdot\mathbf{P}$.
@@
The paper is organized as follows: In the next section, we present experimental studies of domain structures in flat slabs with bidirectional anchoring $\hat{\mathbf{n}}_{0}=-\hat{\mathbf{n}}_{0}$ designed to set uniform planar or radial patterns of molecular orientations at the bounding surfaces. The radial pattern produces a topological defect of charge +1. We observe that the $\mathrm{N}_{\mathrm{F}}$ avoids monocrystal alignment of $\mathbf{P}$ by forming $\pi$-twisted domains ($\pi$-TDs) in planar cells, provided they are sufficiently thick (a few microns). The twist axis is perpendicular to the bounding plates. In thin (micron or less) planar cells, the electrostatic energy is reduced by forming a periodic lattice of elongated “uniform domains” (UDs) with constant $\mathbf{P}$ parallel to $\hat{\mathbf{n}}_{0}$ but alternating the polarity from one domain to the next. In the radial splay patterns, the surface-imposed divergence of $\mathbf{P}$ is relaxed by the TDs in thick cells and by “pie slices” of splay domains (SDs) in thin cells, with the polarization alternatively pointing toward or away from the defect core.
In the theoretical part (Section III), we first review Khachaturyan’s prediction that an unconstrained $\mathrm{N}_{\mathrm{F}}$ has a spontaneously twisted polarization $\mathbf{P}$ with a characteristic period $\lambda_{z}$ in a cylindrical domain and explore how this period depends on the concentration of ions. We also show that the $\mathrm{N}_{\mathrm{F}}$ confined in a planar cell with a bidirectional anchoring will form a $\pi$-twist above a critical cell thickness in order to satisfy the anchoring conditions while decreasing the energetic cost of the depolarization field. In Section IV, we consider polarization domain patterns that occur in thin cells with various pre-patterned anchoring conditions. For bidirectional uniform planar anchoring, $\hat{\mathbf{n}_{0}}=\text{const}$, striped patterns with a characteristic wavelength $\lambda_{x}^{*}$ emerge. In cells with apolar orientation $\mathbf{n}_{0}$ forming a $+1$ radial “aster” defect, the $\mathrm{N}_{\mathrm{F}}$ breaks up into arrays of SDs shaped like “pie slices”. We develop theories for both the stripe width and the number of SD slices. We compare our theoretical predictions and experimental measurements of the number of domain walls (SD slices or stripe widths) in Section V. We draw conclusions and point to future directions in Section VI.
II Experiment
II.1 Materials and methods
We explore an $\mathrm{N}_{\mathrm{F}}$ material abbreviated DIO Nishikawa et al. (2017a) and synthesized in the laboratory as described in Mandle et al. (2017b). On cooling from the isotropic (I) phase, the phase sequence of DIO is I-$174^{\circ}\mathrm{C}$-$82^{\circ}\mathrm{C}$-$\mathrm{SmZ}_{\mathrm{A}}$-$66^{\circ}\mathrm{C}$-$\mathrm{N}_{\mathrm{F}}$-$34^{\circ}\mathrm{C}$-crystal, where $\mathrm{SmZ}_{\mathrm{A}}$ is an antiferroelectric smectic Chen et al. (2023). The sandwich-type cells are bounded by two glass plates with layers of a photosensitive dye Brilliant Yellow (BY), which shows maximum absorption in the range 400 nm to 550 nm. BY is dissolved in dimethylformamide (DMF) at a concentration 0.5 wt %. The filtered BY-DMF solution is spin-coated onto the substrates at 3000 rpm for 30 seconds and baked for 30 minutes at $90^{\circ}\mathrm{C}$. The spin-coating and baking procedures are performed in a humidity-controlled environment with relative humidity fixed at 0.2.
To achieve apolar planar alignment, the BY-coated assembled cell is exposed to a light beam (light source EXFO X-Cite with a spectral range of 320 to 750 nm) with a linear polarizer for 10 minutes. This irradiation induces bidirectional molecular alignment perpendicular to the polarization axis of the normally incident light. The radial aster pattern of the dye molecules at the substrates is induced by irradiating the substrates through a plasmonic metamask with radial arrangements of nanoslits Guo et al. (2016a, b). The bidirectional apolar anchoring is set by the same light beam that passes through the metamask and acquires local linear polarization orthogonal to the long axis of the nanoslit. The BY molecules realign perpendicularly to the local light polarization, thus the pattern of BY molecules replicates the pattern of nanoslits. To ensure that the surface patterns are the same on the top and bottom plates, these plates are assembled into an empty cell with a preset distance $h$ between them and irradiated by the same light beam. The cell is then filled with DIO in the N phase and then cooled down to the $\mathrm{N}_{\mathrm{F}}$ phase.
II.2 Planar cells
@@
Figure 1: Domain structures in photoaligned cells with planar apolar anchoring. (a) A polarizing optical microscopy texture of uniform domains (UDs) in a thin $\mathrm{N}_{\mathrm{F}}$ cell, $h=0.7~{}\mu\mathrm{m}$ (temperature $T=55^{\circ}\mathrm{C}$) with bidirectional anchoring indicated by the white double arrow. (b,c) The same, counterclockwise and clockwise uncrossing of analyzer and polarizer, respectively. Pairs of parallel arrows in (b) illustrate that $\mathbf{P}$ does not change along the $z$-axis normal to the cell but alternates from one UD to the next along the $x$-axis. (d) The same, observation with an optical compensator; the slow axis is along the red double arrow. (e) A polarizing optical microscopy texture of a thick $\mathrm{N}_{\mathrm{F}}$ cell, $h=5~{}\mu\mathrm{m}$ ($T=60^{\circ}\mathrm{C}$) . Light transmission through the cell indicates that the sample is split into $\pi$-twisted domains ($\pi$-TDs), shown schematically by two antiparallel arrows that twist by $\pi$ around the $z$-axis. (f) Domains are suppressed when DIO is doped with 0.5 wt.of an ionic fluid BMIM-$\mathrm{PF}_{\mathrm{6}}$. (g,h,i) The twisted polarization in the $\pi$-TDs is readily recognized by observing the textures with uncrossed (g,i) and crossed (h) polarizers. (j) Fitting the dependence of transmitted light intensity on the angle $\gamma$ between the polarizers yields the opposite twist angles $\tau=\pm 175^{\circ}$ for the two domains highlighted in (h). Textures (e-i) are captured using a green interferometric filter with a center wavelength $\lambda=532~{}\mathrm{nm}$ and bandwidth of $1~{}\mathrm{nm}$.
The planar cells in the N and $\mathrm{SmZ}_{\mathrm{A}}$ phases show homogeneous textures with the molecular director $\hat{\mathbf{n}}$ parallel to the photoinduced apolar “easy axis” $\hat{\mathbf{n}}_{0}=-\hat{\mathbf{n}}_{0}=(0,\pm{1},0)$. Here and in what follows, we use the Cartesian coordinates $(x,y,z)$ in which the $y$-axis is the direction of the anchoring and the $z$-axis is normal to the film.
@@
Upon cooling, the texture remain homogeneous for about $(4-8)~{}{}^{\circ}\mathrm{C}$ below the $\mathrm{SmZ}_{\mathrm{A}}$-$\mathrm{N}_{\mathrm{F}}$ transition point, depending on the cell thickness. Thin cells, $h<2~{}\mu\mathrm{m}$, preserve uniformity, $\mathbf{P}=P(0,\pm{1},0)$ for $(6-8)^{\circ}\mathrm{C}$, after which they split into a lattice of UDs elongated along $\hat{\mathbf{n}}_{0}$, each of width on the order of $10~{}\mu\mathrm{m}$, Fig. 1(a-d). When $\hat{\mathbf{n}}_{0}$ is parallel to one of the polarizers, the UDs are practically extinct between two crossed polarizers, Fig. 1(a), and their optical retardance equals that of the optical compensator when the latter is inserted between the sample and the analyzer, Fig. 1(d). The textures observed with polarizers uncrossed counterclockwise, Fig. 1(b), and clockwise, Fig. 1(c), differ little from each other. One concludes that $\mathbf{P}$ in UDs aligns along $\hat{\mathbf{n}_{0}}$ and their polarity alternates from $\mathbf{P}=P_{0}(0,1,0)$ in one domain to $\mathbf{P}=P_{0}(0,-1,0)$ in the next. There are only few regions, marked with a letter “T” in Figs. 1(b,c), in which the textures with uncrossed polarizers do differ, which suggests a twist of $\mathbf{P}$ along the $z$-axis.
@@
Thick cells, $h>2~{}\mu\mathrm{m}$, show a very different behavior. Below $62^{\circ}\mathrm{C}$, they develop a stripe pattern of $\pi$-twisted domains ($\pi$-TDs), recognized by the absence of light extinction when viewed between two crossed polarizers, one of which is along $\hat{\mathbf{n}}_{0}$, as seen in Fig. 1(e). This $\mathrm{N}_{\mathrm{F}}$ texture is similar to the previously studied TDs with alternating left-handed and right-handed twists in cells in which one plate sets a unipolar alignment of $\mathbf{P}$ and the other is azimuthally degenerate Kumari et al. (2024a).
@@
The addition of an ionic salt 1-Butyl-3-methylimidazolium hexafluorophosphate (BMIM-$\mathrm{PF}_{\mathrm{6}}$) suppresses the $\pi$-TDs, as the sample is predominantly extinct between the crossed polarizers, Fig. 1(f). The added salt also decreases the temperatures of phase transition by approximately $20~{}{}^{\circ}\mathrm{C}$, in agreement with previous studies Zhong et al. (2025); Kumari et al. (2024a); Rupnik et al. (2025). The dependency of a transmitted light intensity on the angle $\gamma$ between the directions of polarization of the polarizer and analyzer allows one to determine that the twist angle $\tau$ between the bottom and top orientation of $\mathbf{P}$ in the DIO cell of a thickness $h=4~{}\mathrm{\mu m}$ is close to $180^{\circ}$, Fig. 1(g-j).
Qualitatively, the observed $\pi$-TDs can be understood as the avoidance of a strong depolarization field at the expense of the twist elastic energy and the energy of the domain walls; the balance of these tendencies and the effect of ions will be discussed in Sections III and IV below.
@@
Note that the domain structures presented in Fig. 1 are different from the recently described splay and double-splay domain structures that form in the material RM734 above the phase transition to the $\mathrm{N}_{\mathrm{F}}$ Rupnik et al. (2025); Ma et al. (2024). These splay domain structures are optically discernable when the material is exposed to an ionic fluid Rupnik et al. (2025) or an ionic polymer Ma et al. (2024). We also observe the splay domain textures in our RM734 samples doped with BMIM-$\mathrm{PF}_{\mathrm{6}}$, but only above the temperature at which the material transitions into the $\mathrm{N}_{\mathrm{F}}$ phase. The splay domains above the $\mathrm{N}_{\mathrm{F}}$ temperature range are attributed to the flexoelectric effect, which favors the splay of molecules with head-tail asymmetry Rupnik et al. (2025); Ma et al. (2024). In the $\mathrm{N}_{\mathrm{F}}$ phase, this flexoelectricity-triggered splay is suppressed by the space charge that accompanies divergence of $\mathbf{P}$ Rupnik et al. (2025). As a result, the domain structures in the $\mathrm{N}_{\mathrm{F}}$ phase, Fig. 1, are different from the domain structures reported in Refs. Rupnik et al. (2025); Ma et al. (2024). Because of this principal difference, in what follows, we consider only the electrostatic, elastic, and surface anchoring effects and disregard the flexoelectric effect.
II.3 Radial patterns
@@
Figure 2: Structure of a +1 radial prepatterned splay defect. (a) A polarizing optical microscopy texture (recorded in a monochromatic light using a green filter with $532~{}\mathrm{nm}$ wavelength and $1~{}\mathrm{nm}$ bandwidth) of an $\mathrm{N}_{\mathrm{F}}$ cell with thickness $h=2~{}\mu\mathrm{m}$, $T=55^{\circ}\mathrm{C}$, and cooling rate $0.1^{\circ}\mathrm{C}$ /min. The parallel arrows show that the polarization at the bottom and top plates are parallel to each other, but alternate from one splay domain (SD) to the next along the azimuthal direction in the $xy$ plane. (b) Similar cell, $h=2~{}\mu\mathrm{m}$, cooling rate $5^{\circ}\mathrm{C}$/min. (c,d) Number of domains as a function of distance $R$ from the defect core for cell thicknesses $h=(1-16)~{}\mu\mathrm{m}$, cooling rates $0.1^{\circ}\mathrm{C}$ and $5^{\circ}\mathrm{C}$, respectively. (e) Number of domains vs $h$ measured at different distances $R$ from the defect core, cooling rate $0.1^{\circ}\mathrm{C}$/min. Filled symbols and solid lines correspond to the total number of domains, while open symbols and dashed lines correspond to $\pi$-TDs. (f) The thickness dependence of the number of domains, cooling rate $5^{\circ}\mathrm{C}$/min. (g) Fraction of TDs as a function of the cell thickness $h$ measured at different distances $R$ from the defect core; cooling rate $0.1^{\circ}\mathrm{C}$ /min. (h) Radial structure with no domains in a thin cell, $h=1~{}\mu\mathrm{m}$, imaged using the Microimager PolScope with the ticks showing the director field $\hat{\mathbf{n}}(x,y)$.

### v2

@@ Abstract
The newly-discovered ferroelectric nematic liquid crystal exhibits a variety of unique defect phenomena. The depolarization field in the material favors spontaneous spatial variations in polarization, manifesting in diverse forms such as bulk twists and arrangements of alternating polarization domains. The configuration of these domains is governed by a balance between depolarization field reduction and molecular alignment at interfaces. We investigate a ferroelectric nematic confined in a thin cell with apolar surface anchoring, patterned using photoalignment. Under uniform planar alignment, the system forms stripes, while a radial +1 defect pattern results in pie-slice domains. Neighboring domains show either opposite directions of uniform polarization (thin cells) or opposite handedness of the spontaneous twist (thick cells). Our calculations and experiments demonstrate that electrostatic interactions tend to shrink domain size, whereas elastic and surface anchoring effects promote larger domains. In this work, we make predictions and measurements of the domain size as a function of cell thickness, and show that ionic screening suppresses domain formation.
keywords:
ferroelectric, nematic liquid crystal, domains, pattern formation
Introduction
@@
The polar ordering of the recently discovered ferroelectric nematic ($\mathrm{N}_{\mathrm{F}}$) liquid crystal [1, 2, 3, 4] creates a fascinating interplay between the elasticity, surface interactions, and the electrostatic energy associated with spatial variations of the spontaneous polarization $\mathbf{P}$. The ensuing $\mathrm{N}_{\mathrm{F}}$ structures are not restricted by crystallographic axes and can be studied by polarizing optical microscopy since, in the materials explored so far, the polarization $\mathbf{P}$ is along the optic axis, or the director $\hat{\mathbf{n}}$.
@@
$\mathrm{N}_{\mathrm{F}}$ structures tend to avoid a splay deformation (divergence of $\mathbf{P}$) since splay creates a bound charge of a bulk density $\rho=-\nabla\cdot\mathbf{P}$ and increases the electrostatic energy. Polarization-related charges are also avoided at surfaces and at domain walls. For example, a uniform polarization, $\mathbf{P}(x,y,z)=\text{const.}$., would deposit charges at the opposite ends of the sample and create a strong, energetically costly depolarization field: $E=-P/\epsilon\epsilon_{0}\approx 10^{8}~{}\mathrm{V}/\mathrm{m}$; here $\epsilon\approx(10-100)$ [5, 6] and $P\approx(3-7)\times 10^{-2}~{}\mathrm{C}/\mathrm{m}^{2}$ are the typical permittivity and polarization of the $\mathrm{N}_{\mathrm{F}}$ phase. Polydomain textures of thin $\mathrm{N}_{\mathrm{F}}$ films that impose no preferred in-plane orientation of $\mathbf{P}$, such as a film supported by an isotropic fluid [7, 8] or freely suspended in air [9], present clear evidence of these tendencies. First, the spontaneous polarization is everywhere in the plane of the film, avoiding surface charges, which would occur whenever $\mathbf{P}$ is tilted. Second, the in-plane textures are dominated by two types of domains, in which $\mathbf{P}$ is either uniform or bends into circular vortices, thus avoiding splay deformations and associated space charge in the bulk. Domain walls separating these domains adopt the shapes of conic sections, such as parabolas and hyperbolas [7, 8, 9].
@@
The $\mathrm{N}_{\mathrm{F}}$ structure changes dramatically when one of the film’s surfaces imposes a unidirectional alignment of $\mathbf{P}$ and the other is azimuthally degenerate. In this case, one might expect a uniform state since circular vortices are not compatible with the unidirectional surface alignment. However, experiments [10] demonstrate that instead of being uniform, the polarization twists around the film normal, so that the vectors $\mathbf{P}$ are antiparallel to each other at the bottom and the top surfaces, thus mitigating the depolarization effect. The twisted structures arise from the balance of electrostatic and elastic energies. Clearly, this balance should be affected by the geometry of confinement (e.g., a film with a large lateral extension or a long cylinder), by surface anchoring at the bounding plates, and by the free ions capable of at least partial screening of bound charges.
@@
In this work, we explore experimentally and theoretically the interplay between electrostatic and elastic energies in $\mathrm{N}_{\mathrm{F}}$ domain structures, taking into account the effects of surface anchoring and ionic screening. The surface anchoring in the experiments is designed to be in-plane apolar (bidirectional) by using a photoalignment technique [11, 12]. There are two reasons: First, apolar anchoring should avoid twists that are artificially created by the antiparallel assembly of two plates with a unidirectional alignment of $\mathbf{P}$, which happens for mechanically rubbed plates. Second, the photoalignment technique allows one to impose various patterns of surface alignment of $\mathbf{P}$ (including an unwelcome splay), thereby exploring the electrostatics-elasticity balance in different contexts.
@@
We find that, generally, the electrostatic interaction prefers a spatial modulation of $\mathbf{P}$ with a characteristic size $\lambda$ (e.g., the twist pitch or the domain size), generating an energetic contribution that increases with $\lambda$. Conversely, any such modulation incurs an elastic or anchoring energy penalty, which decreases with $\lambda$. Thus, the balance between elastic (or anchoring) energy and the electrostatic interaction generates a preferred value of $\lambda$, leading to the various patterns considered here.
@@
We present experimental studies of domain structures in flat slabs with bidirectional anchoring $\hat{\mathbf{n}}_{0}$ designed to set uniform planar or radial patterns of molecular orientations at the bounding surfaces. We observe that the $\mathrm{N}_{\mathrm{F}}$ avoids monocrystal alignment of $\mathbf{P}$ by forming $\pi$-twisted striped domains ($\pi$-TDs) in planar cells, provided they are sufficiently thick (a few microns). The twist axis is perpendicular to the bounding plates. In thin (micron or less) planar cells, the electrostatic energy is reduced by forming a periodic lattice of elongated uniform domains (UDs) with constant $\mathbf{P}$ parallel to $\hat{\mathbf{n}}_{0}$ but alternating the polarity from one domain to the next. In the radial splay patterns, the surface-imposed divergence of $\mathbf{P}$ is relaxed by the TDs in thick cells and by pie slices of splay domains (SDs) in thin cells, with the polarization alternatively pointing toward or away from the $+1$ defect core. We develop theories for both the stripe width and the number of SD slices. We compare our theoretical predictions and experimental measurements of the number of domain walls (SD slices or stripe widths) in cells of varying thickness $h$.
Results
Stripe patterns in planar cells
@@ Results
Figure 1: Domain structures in photoaligned cells with planar apolar anchoring. a A polarizing optical microscopy texture of uniform domains (UDs) in a thin $\mathrm{N}_{\mathrm{F}}$ cell, $h=0.7~{}\mu\mathrm{m}$ (temperature $T=55^{\circ}\mathrm{C}$) with bidirectional anchoring indicated by the white double arrow. b,c The same, counterclockwise and clockwise uncrossing of analyzer and polarizer, respectively. Pairs of parallel arrows in (b) illustrate that $\mathbf{P}$ does not change along the $z$-axis normal to the cell but alternates from one UD to the next along the $x$-axis. d The same, observation with an optical compensator; the slow axis is along the red double arrow. Most domains are uniform in (a-d), with the exception of small areas labeled with a letter T, indicating a $\pi$-twisted region. e Polarizing optical microscopy texture of a thick $\mathrm{N}_{\mathrm{F}}$ cell, $h=5~{}\mu\mathrm{m}$ ($T=60^{\circ}\mathrm{C}$). Light transmission indicates that the sample is split into $\pi$-twisted domains ($\pi$-TDs), shown schematically by two antiparallel arrows that twist by $\pi$ around the $z$-axis. f Domains are suppressed when DIO is doped with 0.5 wt.of an ionic fluid BMIM-$\mathrm{PF}_{\mathrm{6}}$, shown for a cell with $h=4.0~{}\mu\mathrm{m}$ ($T=43^{\circ}\mathrm{C}$). g,h,i The twisted polarization in the $\pi$-TDs is readily recognized by observing the textures with uncrossed (g,i) and crossed (h) polarizers, for a cell with $h=4.0~{}\mu\mathrm{m}$ ($T=60^{\circ}\mathrm{C}$). Textures (e-i) are captured using a green interferometric filter with a center wavelength $\lambda=532~{}\mathrm{nm}$ and $1~{}\mathrm{nm}$ bandwidth. j Theoretical fits (lines) to light transmission data (points) measured through small localized regions in the two domains highlighted in (h) as a function of the angle $\gamma$ between the polarizers. Fits yield twist angles $\tau=\pm 175^{\circ}$ (see Methods) in the two adjacent domains. Error bars are standard deviations of the intensity fluctuations.
@@
The planar cells in the N and $\mathrm{SmZ}_{\mathrm{A}}$ phases show homogeneous textures with the molecular director $\hat{\mathbf{n}}$ parallel to the photoinduced apolar easy axis ±$\pm\hat{\mathbf{n}}_{0}=(0,\pm{1},0)$. Here and in what follows, we use the Cartesian coordinates $(x,y,z)$ in which the $y$-axis is the direction of the anchoring and the $z$-axis is normal to the film. Upon cooling, the texture remain homogeneous for about $(4-8)~{}{}^{\circ}\mathrm{C}$ below the $\mathrm{SmZ}_{\mathrm{A}}$-$\mathrm{N}_{\mathrm{F}}$ transition point, depending on the cell thickness. Thin cells, $h<2~{}\mu\mathrm{m}$, preserve uniformity, $\mathbf{P}=P(0,\pm{1},0)$ for $(6-8)^{\circ}\mathrm{C}$, after which they split into a lattice of UDs elongated along $\hat{\mathbf{n}}_{0}$, each of width on the order of $10~{}\mu\mathrm{m}$, Fig. 1a-d. When $\hat{\mathbf{n}}_{0}$ is parallel to one of the polarizers, the UDs are practically extinct between two crossed polarizers, Fig. 1a, and their optical retardance equals that of the optical compensator when the latter is inserted between the sample and the analyzer, Fig. 1d. The textures observed with polarizers uncrossed counterclockwise, Fig. 1b, and clockwise, Fig. 1c, differ little from each other. One concludes that $\mathbf{P}$ in UDs aligns along $\hat{\mathbf{n}}_{0}$ and their polarity alternates from $\mathbf{P}=P_{0}(0,1,0)$ in one domain to $\mathbf{P}=P_{0}(0,-1,0)$ in the next. There are only few regions, marked with a letter T in Figs. 1b,c, in which the textures with uncrossed polarizers do differ, which suggests a twist of $\mathbf{P}$ along the $z$-axis.
@@
Thick cells, $h>2~{}\mu\mathrm{m}$, show a very different behavior. Below $62^{\circ}\mathrm{C}$, they develop a stripe pattern of $\pi$-twisted domains ($\pi$-TDs), recognized by the absence of light extinction when viewed between two crossed polarizers, one of which is along $\hat{\mathbf{n}}_{0}$, as seen in Fig. 1e. This $\mathrm{N}_{\mathrm{F}}$ texture is similar to the previously studied TDs with alternating left-handed and right-handed twists in cells in which one plate sets a unipolar alignment of $\mathbf{P}$ and the other is azimuthally degenerate [10].
@@
The addition of an ionic salt 1-Butyl-3-methylimidazolium hexafluorophosphate (BMIM-$\mathrm{PF}_{\mathrm{6}}$) suppresses the $\pi$-TDs, as the sample is predominantly extinct between the crossed polarizers, Fig. 1f. The added salt also decreases the temperatures of phase transition by approximately $20{}^{\circ}\mathrm{C}$, in agreement with previous studies [13, 10, 14]. The dependency of a transmitted light intensity on the angle $\gamma$ between the directions of polarization of the polarizer and analyzer allows one to determine that the twist angle $\tau$ between the bottom and top orientation of $\mathbf{P}$ in the DIO cell of a thickness $h=4~{}\mathrm{\mu m}$ is close to $180^{\circ}$, Fig. 1g-j.
@@
Note that the domain structures presented in Fig. 1 are different from the recently described splay and double-splay domain structures that form in the material RM734 above the phase transition to the $\mathrm{N}_{\mathrm{F}}$ [14, 15]. These splay domain structures are optically discernable when the material is exposed to an ionic fluid [14] or an ionic polymer [15]. The splay domains above the $\mathrm{N}_{\mathrm{F}}$ temperature range are attributed to the flexoelectric effect, which favors the splay of molecules with head-tail asymmetry [14, 15]. In the $\mathrm{N}_{\mathrm{F}}$ phase, this flexoelectricity-triggered splay is suppressed by the space charge that accompanies divergence of $\mathbf{P}$ [14]. As a result, the domain structures in the $\mathrm{N}_{\mathrm{F}}$ phase, Fig. 1, are different from the domain structures reported in Refs. [14, 15]. Thus, in what follows, we disregard the flexoelectric effect.
Pie slices in cells pre-patterned with a $+1$ radial splay defect
@@
Figure 2: Characteristics of a +1 radial pre-patterned splay defect. a A polarizing optical microscopy texture (recorded in a monochromatic light using a green filter with $532~{}\mathrm{nm}$ wavelength and $1~{}\mathrm{nm}$ bandwidth) of an $\mathrm{N}_{\mathrm{F}}$ cell with thickness $h=2~{}\mu\mathrm{m}$, $T=55^{\circ}\mathrm{C}$, and cooling rate $0.1^{\circ}\mathrm{C}$ /min. The parallel arrows show the parallel polarizations at the bottom and top plates, with $\pi$ flips from one splay domain (SD) to the next along the azimuthal direction. b Similar cell but recorded under white light, $h=2~{}\mu\mathrm{m}$, $T=55^{\circ}\mathrm{C}$, and cooling rate $5^{\circ}\mathrm{C}$/min. c,d Number of domains as a function of distance $R$ from the defect core for cell thicknesses $h=(1-16)~{}\mu\mathrm{m}$, cooling rates $0.1^{\circ}\mathrm{C}$ and $5^{\circ}\mathrm{C}$, respectively. e Number of domains versus $h$ measured at different distances $R$ from the defect core, cooling rate $0.1^{\circ}\mathrm{C}$/min. Filled symbols and solid lines correspond to the total domain count, while open symbols and dashed lines correspond to the number of $\pi$-TDs. f The $h$-dependence of the number of domains, cooling rate $5^{\circ}\mathrm{C}$/min. Error bars in (c-f) are estimates of uncertainties in the domain count in the sample. g Fraction of TDs as a function of the cell thickness $h$ measured at different distances $R$ from the defect core; cooling rate $0.1^{\circ}\mathrm{C}$ /min. h Radial structure with no domains in a thin cell, $h=1~{}\mu\mathrm{m}$ ($T=50{}^{\circ}\mathrm{C}$), imaged using the Microimager PolScope with the ticks showing the director field $\hat{\mathbf{n}}(x,y)$. The thin blue line of decreased retardance is a domain wall separating the $+1/2$ defects that split from the central $+1$ splay pattern.